OpenStudy (timmyk14):

Can someone let me know whether or not I'm correct? Please, Medal+Fan. 2. Describe how you would graph this line using the slope-intercept method. Be sure to write in complete sentences. I would plot the ordered pair (0,400) on the coordinate plane. Then to find the other ordered pair you would simply use the slope (-2,3). So, go two spaces to the right of (0,400) and then three up from (-2,400). The ordered pair is (-2,397).

OpenStudy (timmyk14):

1. Change the equation into slope-intercept form. Identify the slope and y-intercept of the equation. Be sure to show all of your work. 2x + 3y = 1200 2x-1200=-3y 1/3*(2x-1200)=1/3*-3y 2/3x-400=-y -2/3x+400=y

OpenStudy (anonymous):

That is correct, although, with slope, verticality comes first. Good job.

OpenStudy (anonymous):

Oh, and y comes first on slope-int. form